# Towards the new XAI: A Hypothesis-Driven Approach to Decision Support Using Evidence

### v2

@@ Abstract.
XAI, decision-making, hypothesis-driven, evidence, evaluative AI
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Artificial intelligenceccs: Human-centered computing HCI theory, concepts and models
1. Introduction
@@
Research has shown that AI recommendations, even when accompanied with explanations, are not always helpful in supporting decision-making (Bansal et al., 2021; Wang and Yin, 2022; Buçinca et al., 2021; Gajos and Mamykina, 2022). The direct causes of this are under-reliance and over-reliance (Vered et al., 2023). With under-reliance, decision-makers reject AI recommendations, even when they may be correct. Alternatively, decision-makers may overly rely on AI recommendations, hence be led to errors when the AI is incorrect. In either case, they tend to fixate on a particular hypothesis without sufficiently considering others (Miller, 2023). Approaches such as cognitive forcing, based on ideas from human psychology, have been proposed to address limitations of the AI recommendation approach (Buçinca et al., 2021), with recent work indicting that withholding AI model the recommendations, at least for a short time, while still providing the user with an explanation of that recommendation, can be helpful (Gajos and Mamykina, 2022). Recently, Miller (2023) proposed a so-called hypothesis-driven decision-making paradigm called evaluative AI. The main aim of this paradigm is to focus the decision loop on the human decision maker, providing them with the right evidence to support their own intuitions, rather than focusing the decision loop on machine recommendations. This paradigm offers a promising direction in building better decision support in explainable AI (XAI) research by focusing on human decision makers considering multiple possible hypotheses.
@@
In this paper, we describe and evaluate an approach for building a hypothesis-driven decision-making model that uses the Weight of Evidence (WoE) framework (Alvarez Melis et al., 2021a). To the best of our knowledge, this is the first work that compares empirically and in a controlled manner the hypothesis-driven approach (Miller, 2023) with two other popular decision-making approaches (recommendation-driven and AI-explanation-only (Gajos and Mamykina, 2022)). Our contributions are:
@@
  • The Evidence-Informed Hypothesis-Driven Decision Making model, building on the Weight of Evidence (WoE) framework to the hypothesis-driven approach.
@@
  • Two human behavioural experiments comparing our hypothesis-driven approach with two common decision-aid approaches: (1) the standard model recommendation with explanation; and (2) a form of cognitive forcing that provides only AI explanations (Gajos and Mamykina, 2022). The results show that hypothesis-driven approach improved decision accuracy and reduces over-reliance compared to standard recommendation-driven approaches, at the cost of a slight increase in under-reliance. Furthermore, the hypothesis-driven reducesd under-reliance significantly compared to the AI-explanation-only. Our qualitative analysis identifies some limitations and challenges in the three approaches, and shows that participants used the hypothesis-driven approach in a materially different way than the recommendation-driven or AI-explanation-only conditions, with participants focusing more on the evidence than on their own background knowledge.
2. Background and Related Work
@@
In this section, we give a brief overview of related work on common AI-assisted decision-making paradigms. We then highlight key literature about applying explainable AI in supporting decision-making.
2.1. AI-Assisted Decision Making Paradigms
@@
In the AI-first decision-making workflow, it has been demonstrated that participants feel more confident and are also faster in decision-making (Fogliato et al., 2022). These participants also rated AI as more practical. In terms of limitations, the anchoring effect is reported to occur more often in the AI-first workflow (Fogliato et al., 2022; Rastogi et al., 2022; Buçinca et al., 2021) in which people overly rely on the AI recommendation (also called over-reliance). Anchoring effect (Tversky and Kahneman, 1974) refers to giving stronger preference to the earlier knowledge rather than doing a full revision and considering the latest evidence. By contrast, Fogliato et al. (2021) did not find any significant difference in the participants’ performance, which is measured by accuracy between the two workflows (AI-first and human-first). However, they also found that participants are 65% more likely to revise their answers in the human-first setting than those in the AI-first setting.
@@
Human-first decision workflow has been shown to help reduce the reliance on erroneous AI recommendations (Buçinca et al., 2021; Fogliato et al., 2022). However, experts may interact with decision-making systems differently from laypeople (crowdworkers). For example, Fogliato et al. (2022); Gaube et al. (2021) ran studies with radiologists who were the experts. Their task is to review patients’ X-ray images. The studies conclude that in human-first workflow with expert participants, they are less likely to leverage AI advice even though the AI is more accurate. In fact, this is referred to algorithm aversion (Dietvorst et al., 2015) or under-reliance.
A human-first paradigm called cognitive forcing, based on earlier ideas in psychology for interventions that elicit human thinking at decision-making time (Lambe et al., 2016), has been proposed as a way to improve users’ engagement and also increase their learning when interacting with the AI (Buçinca et al., 2021; Gajos and Mamykina, 2022). Four cognitive forcing designs have been introduced: (1) On demand: Participants can only see the AI recommendation when they request it; (2) Update: Participants first made a decision without seeing the AI recommendation. Then, they were shown the AI prediction and could update their decision later; (3) Wait: Participants had to wait for 30 seconds before the AI decision was shown; and (4) Only AI explanation: Providing just the AI explanation and no AI recommendation, on the basis that this may help people process the AI explanation more carefully and therefore, improve their knowledge and make better decisions (Gajos and Mamykina, 2022). Importantly, cognitive forcing has been shown to reduce over-reliance compared to the standard AI suggestion approach, although that study had a limitation in that the AI prediction was always correct. There are also some recognised trade-offs of cognitive forcing designs: more time-consuming (Gajos and Mamykina, 2022), and less trust (Buçinca et al., 2021).
@@
Miller (2023) argues that AI-assisted decision support is on the cusp of a paradigm shift. This shift is away from the idea of human-first or AI-first, and into a framework he calls evaluative AI. The key insight of evaluative AI is to not necessarily provide a recommendation, and instead to support the human cognitive decision-making process by providing evidence that supports or against the particular hypothesis that a human decision maker is considering. Miller argues that this would help to prevent over- and under-reliance, and would help the decision maker to retain their internal locus on control (Shneiderman et al., 2016).
2.2. Explainable AI (XAI) in Supporting Decision Making
@@
There is no straightforward position regarding when humans are well calibrated to accept AI-generated advice (Vered et al., 2023). Overall, study participants appear more likely to accept an AI’s recommendation when provided with explanations, regardless of the model’s correctness (Bansal et al., 2021; Jacobs et al., 2021; van der Waa et al., 2021). Explanations can increase the accuracy of the human-AI team when the AI is correct, but decrease it when it is wrong, resulting in over-reliance on the AI’s recommendations. Arguably, this is because the current explanation styles do not provide details of the underlying rationale of the AI model behaviour (van der Waa et al., 2021). We therefore should be careful when selecting the explanation type as it can have a significant effect on whether users decide to rely on them  (Caruana et al., 2015; Lombrozo, 2007).
@@
In this paper, we will generate evidence-based explanations, which are similar to feature importance explanations. The main difference is that Weight of Evidence uses log likelihoods and log odds ratios to generate explanations, whereas LIME (Ribeiro et al., 2016) and SHAP (Lundberg and Lee, 2017) find feature importance by modifying the predictive posterior probability in various ways.
Evidence-based explanations have been applied to support decision-making and debug models in several prior research (Kulesza et al., 2011, 2015; Poulin et al., 2006) other than from Alvarez Melis et al. (2021a). A closely related work to Alvarez Melis et al. (2021a) is from Poulin et al. (2006). Poulin et al. (2006) propose a framework called ExplainD that uses additive evidence. The framework also measures the weight of evidence using a Naive Bayes classifier along with highlighting the negative and positive evidence for a decision. However, the problem being considered is a binary classification. Furthermore, there is still room for improvement by conducting experiments to evaluate the framework. Kulesza et al. (2011, 2015) introduce EluciDebug in email classification using Multinomial Naive Bayes classifier (MNB). The EluciDebug prototype provides an interface that includes important words and the folder size that both contribute to the email classification. However, the prototype did not specifically give positive and negative evidence in decision-making situations.
3. Evidence-Informed Hypothesis-Driven Decision Making Model
We define the evidence-informed hypothesis-driven decision-making model by implementing the evaluate AI (hypothesis-driven) paradigm (Miller, 2023) using the WoE model. Specifically, given a classification problem, decision makers explore evidence for and against each hypothesis (i.e. an output class). We allow decision-makers to interact with the model by repeatedly selecting a hypothesis for which they can then see the positive (or negative) evidence.
@@ 3. Evidence-Informed Hypothesis-Driven Decision Making Model
We generate the weight of evidence for possible hypotheses using Weight of evidence (WoE), which is a probabilistic approach for analysing variable importance, introduced in the context of explainability by Alvarez Melis et al. (2021a) building on Good (1985). It provides a quantitative response to the question of why a model predicted output $h$ for a particular input $x$ in terms of how much each input feature $x_{i}$ provides in favour of, or against, $h$, relative to alternatives. Through Bayes rule, WoE can be understood as an adjustment to the prior log odds caused by observing the evidence.
@@
(1) $\operatorname*{woe}(h\mid x_{i})=\log\frac{P(x_{i}\mid h)}{P(x_{i}\mid\bar{h})}=\log\frac{P(h\mid x_{i})}{P(\bar{h}\mid x_{i})}-\log\frac{P(h)}{P(\bar{h})}$
@@
To assist users with interpretability, Alvarez Melis et al. (2021a) complement the display of the magnitude of the weight of the evidence with a notion of significance level of the evidence, using a scale of seven categories: decisive-against, strong-against, substantial-against, not-significant, substantial-in-favor, strong-in-favor, decisive-in-favor. The details can be found in the rule-of-thumb guidelines here (Alvarez Melis et al., 2021b).
@@
Formally, by considering the importance of the evidence, we re-define the total weight of evidence from a human decision making perspective as follows:
(4) $\operatorname*{woe}(h)=\sum_{i=1}^{n}\gamma_{i}\times\operatorname*{woe}(h\mid x_{i})$
@@ 4. Experiment Design
In this section, we describe the task implemented in the human behaviour experiment and the experiment design. In selecting a decision-making task, we identified requirements similar to those used in other studies of how explanations can assist human decision makers interacting with AI decision support, e.g. Vasconcelos et al. (2023): the task should not be too easy for humans to complete without a decision aid, but also, as we were using lay subjects from Prolific for this particular study, the task cannot require specialist knowledge.
@@
We chose a version of the housing price prediction task studied previously in an XAI context (Abdul et al., 2020; Chiang and Yin, 2022). In this task, participants are provided with information about house features and with other information which varies by experimental condition, and are asked to choose whether the given house would have a sale price of low, medium or high. As noted by others (Chiang and Yin, 2022), real estate valuation is a domain where ML models have been developed to help people make better decisions, predicting house prices is a task that lay people may need to do in real life, so it is not unrealistic to expect they have sufficient day-to-day knowledge to make predictions and decide whether or not to rely on an AI model.
@@
Experimenting with this task, we compared the hypothesis-driven approach with two state-of-the-art decision-making approaches using quantitative measures for efficiency, performance and reliance and a qualitative analysis of information use. In the terminology of a recent review of XAI evaluation (Lai et al., 2023), the first two points of comparison are a form of evaluation with respect to the decision task, and the latter two focus on users’ perception and use of the AI system itself.
4.1. Dataset and Model Implementation
@@
To build our model, we used the Ames Housing Dataset (De Cock, 2011) and the open source code on GitHub (Tan, 2021) for data pre-processing. The data after pre-processing has a total of 2616 instances and 28 features. We processed the dataset further by converting the house price into three output classes (low price, medium price and high price). We also balanced the dataset to ensure that the three classes had the same number of instances by using Near-Miss Undersampling. Finally, we had a total of 1920 instances with 640 instances for each class.
@@
We selected six features for the human experiment in the house-price decision making task by applying Gradient Boosting Classification model over the data. Considering domain specific decision making about house prices, we propose there to be three important features (quality of construction, house age and location) and three unimportant features (fireplaces, kitchen quality and central air conditioning). We divided the dataset into 80% for the training set and 20% for the test set. Following Alvarez Melis et al. (2021a), we use a Gaussian Naïve Bayes (GNB) classifier to obtain $P(x_{i}\mid h)$. This assumes that features are independent, but the model and implementation work for any probabilistic classifier. We chose this model because it is a simple discriminative classifier that aligns with previous work on evidence-based explanations (Poulin et al., 2006; Kulesza et al., 2015).
4.2. Experimental Conditions
@@
Although participants in the AI-explanation-only and hypothesis-driven conditions did not see a recommendation, it was expected that the displayed information from the WoE framework would provide insight that participants could use to support their decision making. We note a similar AI-explanation-only approach has been explored previously (Gajos and Mamykina, 2022).
4.3. Research Questions and Hypotheses

### v3

@@ Abstract
[…]
496
1 Introduction
@@
Research has shown that AI recommendations, even when accompanied with explanations, are not always helpful in supporting decision-making [4, 39, 6, 15]. The direct causes of this are under-reliance and over-reliance [38]. With under-reliance, decision-makers reject AI recommendations, even when they may be correct. Alternatively, decision-makers may overly rely on AI recommendations, hence be led to errors when the AI is incorrect. In either case, they tend to fixate on a particular hypothesis without sufficiently considering others [28]. Approaches such as cognitive forcing, based on ideas from human psychology, have been proposed to address limitations of the AI recommendation approach [6], with recent work indicting that withholding AI model the recommendations, at least for a short time, while still providing the user with an explanation of that recommendation, can be helpful [15]. Recently, Miller [28] proposed a so-called hypothesis-driven decision-making paradigm called evaluative AI. The main aim of this paradigm is to focus the decision loop on the human decision maker, providing them with the right evidence to support their own intuitions, rather than focusing the decision loop on machine recommendations. This paradigm offers a promising direction in building better decision support in explainable AI (XAI) research by focusing on human decision makers considering multiple possible hypotheses.
@@
In this paper, we describe and evaluate an approach for building a hypothesis-driven decision-making model that uses the Weight of Evidence (WoE) framework [2]. To the best of our knowledge, this is the first work that compares empirically and in a controlled manner the hypothesis-driven approach [28] with two other popular decision-making approaches (recommendation-driven and AI-explanation-only [15]). Our contributions are:
@@
  • The Evidence-Informed Hypothesis-Driven Decision-Making model, building on the Weight of Evidence (WoE) framework to the hypothesis-driven approach;
@@
  • Two human behavioural experiments comparing our hypothesis-driven approach with two common decision-aid approaches: (1) the standard model recommendation with explanation; and (2) a form of cognitive forcing that provides only AI explanations [15]. The results show that hypothesis-driven approach improved decision accuracy and reduces over-reliance compared to standard recommendation-driven approaches, at the cost of a slight increase in under-reliance. Furthermore, the hypothesis-driven approach reduces under-reliance significantly compared to the AI-explanation-only approach. Our qualitative analysis identifies some limitations and challenges in the three approaches, and shows that participants used the hypothesis-driven approach in a materially different way than the recommendation-driven or AI-explanation-only conditions, with participants focusing more on the evidence than on their own background knowledge.
2 Background and Related Work
@@
In this section, we give a brief overview of related work on common AI-assisted decision-making paradigms. We then highlight key literature on applying explainable AI in supporting decision-making.
2.1 AI-Assisted Decision-Making Paradigms
@@
In the AI-first decision-making workflow, it has been demonstrated that participants feel more confident and are also faster in decision-making [14]. These participants also rated AI as more practical. In terms of limitations, the anchoring effect is reported to occur more often in the AI-first workflow [14, 30, 6] in which people overly rely on the AI recommendation (also called over-reliance). The anchoring effect [35] refers to giving stronger preference to the earlier knowledge rather than doing a full revision and considering the latest evidence. By contrast, Fogliato et al. [13] did not find any significant difference in the participants’ performance, which is measured by accuracy between the two workflows (AI-first and human-first). However, they also found that participants are 65% more likely to revise their answers in the human-first setting than those in the AI-first setting.
@@
Human-first decision workflow has been shown to help reduce the over-reliance on erroneous AI recommendations [6, 14]. However, experts may interact with decision-making systems differently from laypeople (crowdworkers). For example, Fogliato et al. [14], Gaube et al. [16] run studies with radiologists who were the experts. Their task is to review patients’ X-ray images. The studies conclude that in human-first workflow with expert participants, they are less likely to leverage AI advice even though the AI is more accurate. In fact, this is referred to algorithm aversion [12] or under-reliance.
A human-first paradigm called cognitive forcing, based on earlier ideas in psychology for interventions that elicit human thinking at decision-making time [23], has been proposed as a way to improve users’ engagement and also increase their learning when interacting with the AI [6, 15]. Four cognitive forcing designs have been introduced: (1) On demand: Participants can only see the AI recommendation when they request it; (2) Update: Participants first made a decision without seeing the AI recommendation. Then, they were shown the AI prediction and could update their decision later; (3) Wait: Participants had to wait for 30 seconds before the AI decision was shown; and (4) Only AI explanation: Providing just the AI explanation and no AI recommendation, on the basis that this may help people process the AI explanation more carefully and therefore, improve their knowledge and make better decisions [15]. Importantly, cognitive forcing has been shown to reduce over-reliance compared to the standard AI suggestion approach, although that study had a limitation in that the AI prediction was always correct. There are also some recognised trade-offs of cognitive forcing designs: more time-consuming [15], and less trust [6].
@@
Miller [28] argues that AI-assisted decision support is on the cusp of a paradigm shift. This shift is away from the idea of human-first or AI-first, and into a framework he calls evaluative AI. The key insight of evaluative AI is to not necessarily provide a recommendation, and instead to support the human cognitive decision-making process by providing evidence for or against the particular hypothesis that a human decision maker is considering. Miller argues that this would help to prevent over- and under-reliance, and would help the decision maker to retain their internal locus on control [32].
2.2 Explainable AI (XAI) in Decision Support
@@
There is no straightforward position regarding when humans are well calibrated to accept AI-generated advice [38]. Overall, study participants appear more likely to accept an AI’s recommendation when provided with explanations, regardless of the model’s correctness [4, 19, 36]. Explanations can increase the accuracy of the human-AI team when the AI is correct, but decrease it when it is wrong, resulting in over-reliance on the AI’s recommendations. Arguably, this is because the current explanation styles do not provide details of the underlying rationale of the AI model behaviour [36]. We therefore should be careful when selecting the explanation type as it can have a significant effect on whether users decide to rely on them  [7, 26].
@@
In this paper, we will generate evidence-based explanations, which are similar to feature importance explanations. The main difference is that Weight of Evidence uses log likelihoods and log odds ratios to generate explanations, whereas LIME [31] and SHAP [27] find feature importance by modifying the predictive posterior probability in various ways.
Evidence-based explanations have been applied to support decision-making and debug models in several prior research [20, 21, 29] other than from Alvarez Melis et al. [2]. A closely related work to Alvarez Melis et al. [2] is from Poulin et al. [29]. Poulin et al. [29] propose a framework called ExplainD that uses additive evidence. The framework also measures the weight of evidence using a Naive Bayes classifier along with highlighting the negative and positive evidence for a decision. However, the problem being considered is a binary classification. Furthermore, there is still room for improvement by conducting experiments to evaluate the framework. Kulesza et al. [20, 21] introduce EluciDebug in email classification using Multinomial Naive Bayes classifier (MNB). The EluciDebug prototype provides an interface that includes important words and the folder size that both contribute to the email classification. However, the prototype did not specifically give positive and negative evidence in decision-making situations.
3 Evidence-Informed Hypothesis-Driven Decision-Making Model
We define the evidence-informed hypothesis-driven decision-making model by implementing the evaluative AI (hypothesis-driven) paradigm [28] using the WoE model. Specifically, given a classification problem, decision makers explore evidence for and against each hypothesis (i.e. an output class). We allow decision-makers to interact with the model by repeatedly selecting a hypothesis for which they can then see the positive (or negative) evidence.
@@ 3 Evidence-Informed Hypothesis-Driven Decision-Making Model
We generate the weight of evidence for possible hypotheses using Weight of evidence (WoE), which is a probabilistic approach for analysing variable importance, introduced in the context of explainability by Alvarez Melis et al. [2] building on the approach of Good [17]. It provides a quantitative response to the question of why a model predicted output $h$ for a particular input $x$ in terms of how much each input feature $x_{i}$ provides in favour of, or against, $h$, relative to alternatives. Through Bayes rule, WoE can be understood as an adjustment to the prior log odds caused by observing the evidence.
@@
$\operatorname*{woe}(h\mid x_{i})=\log\frac{P(x_{i}\mid h)}{P(x_{i}\mid\bar{h})}=\log\frac{P(h\mid x_{i})}{P(\bar{h}\mid x_{i})}-\log\frac{P(h)}{P(\bar{h})}$ (1)
@@
To assist users with interpretability, Alvarez Melis et al. [2] complement the display of the magnitude of the weight of the evidence with a notion of significance level of the evidence, using a scale of seven categories: decisive-against (), strong-against (), substantial-against (-), not-significant (N), substantial-in-favor (+), strong-in-favor (++), decisive-in-favor ($+$). The details can be found in the rule-of-thumb guidelines here [3].
@@
Formally, by considering the importance of the evidence, we re-define the total weight of evidence from a human decision-making perspective as follows:
$\operatorname*{woe}(h)=\sum_{i=1}^{n}\gamma_{i}\times\operatorname*{woe}(h\mid x_{i})$ (4)
@@ 4 Experiment Design
In this section, we describe the task implemented in the human behaviour experiment and the experiment design. In selecting a decision-making task, we identified requirements similar to those used in other studies of how explanations can assist human decision makers interacting with AI decision support, e.g. Vasconcelos et al. [37]: the task should not be too easy for humans to complete without a decision aid, but also, as we were using lay subjects from Prolific for this particular study, the task cannot require specialist knowledge.
@@
We chose a version of the housing price prediction task studied previously in an XAI context [1, 10]. In this task, participants are provided with information about house features and with other information which varies by experimental condition, and are asked to choose whether the given house would have a sale price of low, medium or high. As noted by others [10], real estate valuation is a domain where ML models have been developed to help people make better decisions, predicting house prices is a task that lay people may need to do in real life, so it is not unrealistic to expect they have sufficient day-to-day knowledge to make predictions and decide whether or not to rely on an AI model.
@@
Experimenting with this task, we compared the hypothesis-driven approach with two state-of-the-art decision-making approaches using quantitative measures for efficiency, performance and reliance and a qualitative analysis of information use. In the terminology of a recent review of XAI evaluation [22], the first two points of comparison are a form of evaluation with respect to the decision task, and the latter two focus on users’ perception and use of the AI system itself.
4.1 Dataset and Model Implementation
@@
To build our model, we used the Ames Housing Dataset [11] and the open source code on GitHub [34] for data pre-processing. The data after pre-processing has a total of 2616 instances and 28 features. We processed the dataset further by converting the house price into three output classes (low price, medium price and high price). We also balanced the dataset to ensure that the three classes had the same number of instances by using Near-Miss Undersampling. Finally, we had a total of 1920 instances with 640 instances for each class.
@@
We selected six features for the human experiment in the house-price decision-making task by applying Gradient Boosting Classification model over the data. Considering domain specific decision-making about house prices, we propose there to be three important features (quality of construction, house age and location) and three unimportant features (fireplaces, kitchen quality and central air conditioning). We divided the dataset into 80% for the training set and 20% for the test set. Following Alvarez Melis et al. [2], we use a Gaussian Naïve Bayes (GNB) classifier to obtain $P(x_{i}\mid h)$. This assumes that features are independent, but the model and implementation work for any probabilistic classifier. We chose this model because it is a simple discriminative classifier that aligns with previous work on evidence-based explanations [29, 21].
4.2 Experimental Conditions
@@
Although participants in the AI-explanation-only and hypothesis-driven conditions did not see a recommendation, it was expected that the displayed information from the WoE framework would provide insight that participants could use to support their decision-making. We note a similar AI-explanation-only approach has been explored previously [15].
4.3 Research Questions and Hypotheses